# cadrille: Multi-modal CAD Reconstruction with Online Reinforcement Learning

### v2

@@ Abstract
Computer-Aided Design (CAD) plays a central role in engineering and manufacturing, making it possible to create precise and editable 3D models. Using a variety of sensor or user-provided data as inputs for CAD reconstruction can democratize access to design applications. However, existing methods typically focus on a single input modality, such as point clouds, images, or text, which limits their generalizability and robustness. Leveraging recent advances in vision-language models (VLM), we propose a multi-modal CAD reconstruction model that simultaneously processes all three input modalities. Inspired by large language model (LLM) training paradigms, we adopt a two-stage pipeline: supervised fine-tuning (SFT) on large-scale procedurally generated data, followed by reinforcement learning (RL) fine-tuning using online feedback, obtained programatically. Furthermore, we are the first to explore RL fine-tuning of LLMs for CAD tasks demonstrating that online RL algorithms such as Group Relative Preference Optimization (GRPO) outperform offline alternatives. In the DeepCAD benchmark, our SFT model outperforms existing single-modal approaches in all three input modalities simultaneously. More importantly, after RL fine-tuning, cadrille sets new state-of-the-art on three challenging datasets, including a real-world one.
@@ 1 Introduction
Computer-Aided Design (CAD) is the core of modern engineering and manufacturing, providing the tools to create detailed and modifiable 3D models briere2012comparing-3d-cad . Creating CAD models manually requires skills, time, and effort. To simplify this process, CAD reconstruction aims at generating CAD models directly from scanned objects, making the process faster, cheaper, and more accessible overall rukhovich2024cad-recode .
@@
Typically, CAD models are created with a sequence of 2D sketches and 3D operations willis2021fusion360 ; wu2021deepcad . This representation allows CAD models to be easily edited, making it prevalent in popular CAD tools like SolidWorks and AutoCAD and in CAD generation research. Most existing CAD generation methods define CAD sequences using special command tokens khan2024cad-signet ; wu2021deepcad . However, state-of-the-art results are obtained via mapping CAD sequences to casual Python code rukhovich2024cad-recode . Following the same paradigm, we generate CAD models as executable Python scripts.
@@
The most well-studied input modality in CAD reconstruction is naturally a point cloud rukhovich2024cad-recode . However, point clouds can only be obtained when a physical 3D object is available, while scanning usually requires special equipment, making the process complicated for non-experts. Images capture finer details and can be sourced using customer low-end devices (e.g., smartphone cameras), hence relaxing the hardware requirements chen2025cadcrafter ; chen2024img2cad . In the meantime, textual descriptions can enrich the object representation with semantic context khan2024text2cad . Using various input modalities, such as multi-view images, or natural language descriptions, would make design assistance applications simple even for non-experienced users. However, existing approaches typically focus on a single modality at a time, limiting their robustness and generality.
@@
Figure 1: Compared to CAD-Recode, the only existing method of CAD reconstruction as LLM-based code generation, cadrille has two key novelties. First, it goes beyond the standard training scheme and pioneers LLM RL fine-tuning for CAD reconstruction (left). Moreover, besides point clouds only accepted by single-modal CAD-Recode, cadrille extends to images and textual descriptions, making it the first multi-modal approach delivering state-of-art results (right).
The task of Python code generation given various inputs naturally refers to the large Vision-Language models (VLM), which demonstrate strong reasoning capabilities across diverse modalities shao2024deepseekmath ; wang2024qwen2-vl ; yang2024qwen2 . In this work, we harness the power of VLMs to build a multimodal CAD reconstruction model that simultaneously processes point clouds, multi-view images, and texts, and generates Python code.
@@
Existing CAD reconstruction methods face generalization issues due to how they are trained. Specifically, handcrafted CAD datasets are small and limited in diversity, while models trained with procedurally generated data struggle to transfer to the real-world domain khan2024cad-signet ; wu2021deepcad . Inspired by the standard LLM training pipelines shao2024deepseekmath , we introduce a multi-stage training paradigm in the context of CAD reconstruction. We use voluminous procedurally generated data for supervised training, while valuable but scarcer handcrafted data is reserved for RL fine-tuning. This scheme eliminates the need for large-scale handcrafted data and allows the model to first generalize across the CAD domain and then specialize using preference-based objectives. We test Direct Preference Optimization (DPO) rafailov2023dpo , which was employed for the CAD reconstruction in  chen2025cadcrafter . Moreover, since online RL techniques dominate in tasks where feedback can be programatically computed, we also try the online Group Relative Preference Optimization (GRPO) technique. Our experiments demonstrate that online RL is more effective for CAD reconstruction, which aligns with the results obtained in other domains shao2024deepseekmath .
@@
Our experiments show that cadrille  outperforms existing modality-specific baselines in accuracy. Moreover, RL fine-tuning ensures validity of generated Python code, which posed a challenge for prior works. As a result, the proposed approach demonstrates unattainable robustness and sets a new state-of-the-art on several CAD datasets, including a real-world CC3D mallis2023cc3d . Essentially, this opens up new possibilities for generalization in open-world scenarios.
@@
  • We propose cadrille, a CAD reconstruction model that operates jointly on point cloud, images, and text modalities;
  • We are the first to explore RL fine-tuning of LLM models in the context of CAD reconstruction. Specifically, we propose a novel paradigm to split procedurally generated and handcrafted data between supervised fine-tuning (SFT) and RL fine-tuning stages;
  • With a single model, we simultaneously achieve state-of-the-art reconstruction quality for all input modalities on the DeepCAD dataset. RL fine-tuning pushes the quality even further in three benchmarks, including real-world CC3D.
@@ 2 Related Work
Existing CAD generation methods can be classified into three categories based on CAD model representations: constructive solid geometry (CSG)  du2018inversecsg ; ellis2019write ; friedrich2019optimizing ; kania2020ucsg-net ; nandi2018functional ; ren2021csg-stump ; sharma2018csgnet ; tian2019learning ; yu2023d2csg ; yu2022capri-net , boundary representation (B-rep) guo2022complexgen ; jayaraman2022solidgen ; lambourne2021brepnet ; li2019supervised ; li2025caddreamer ; liu2024nvdnet ; liu2024point2cad ; sharma2020parsenet ; smirnov2021learning ; wang2022neural ; wang2020pie-net ; xu2024brepgen and CAD sequence badagabettu2024query2cad ; chen2024img2cad ; chen2025cadcrafter ; dupont2024transcad ; khan2024cad-signet ; khan2024text2cad ; lambourne2022prismcad ; ma2024cad-diffuser ; mallis2024cad-assistant ; ren2022extrudenet ; rukhovich2024cad-recode ; wang2025cad-gpt ; wu2021deepcad ; xu2022skexgen ; xu2023hnc-cad ; yuan2025cad-editor ; zhang2024flexcad . In the CSG foley1996csg paradigm, CAD is represented as a CSG tree constructed using boolean operations (union, subtraction, difference) of geometric primitives (e.g., cubes, cylinders, or spheres). This approach fails to express intricate shapes and is generally not well-aligned with how engineers and designers actually build CAD models. B-rep ansaldi1985brep is a graph that describes connections between faces, edges, and vertices of a 3D model. Creating a B-Rep requires enforcing topological consistency on edges, which introduces additional complexity to the generation procedure and complicates editing of generated models.
@@
Point clouds are the most well-studied input modality in CAD reconstruction. The seminal work on point cloud-based CAD reconstruction by Wu et al. wu2021deepcad proposed encoding CAD sketch-and-extrude sequences as special tokens. Beyond that, DeepCAD, a large-scale dataset of 180k hand-crafted CAD models, was presented. Subsequent works dupont2024transcad ; khan2024cad-signet ; xu2023hnc-cad adopted the same CAD representation and trained on the same DeepCAD dataset. More recently, CAD-Recode rukhovich2024cad-recode introduced a paradigm shift by representing CAD models as Python code, providing greater expressiveness and flexibility, and released a new training dataset of approximately 1 million procedurally generated CAD samples.
@@
Only few recent works chen2025cadcrafter ; chen2024img2cad ; khan2024text2cad ; wang2025cad2program have explored CAD reconstruction from alternative input modalities, such as single- or multi-view images and natural language descriptions. These approaches extend the DeepCAD dataset by rendering synthetic views or generating textual captions for existing CAD models. Among them, CADCrafter chen2025cadcrafter stands out for its unified framework that handles both single- and multi-view inputs, whether rendered or real. In the domain of text-to-CAD, the most notable results come from Text2CAD khan2024text2cad , which uses a vision-language model (VLM) to generate detailed captions for DeepCAD shapes and then trains a model to predict the corresponding sketch-and-extrude sequences from these descriptions.
@@
Generally, state-of-the-art CAD reconstruction approaches are tailored to process specific input modalities with distinct architectures, while multimodal CAD reconstruction remains underexplored. Recent CAD-GPT wang2025cad-gpt reconstructs a CAD model given a single image and textual description, yet falls behind CADCrafter chen2025cadcrafter and Text2CAD khan2024text2cad by a large margin (up to two orders of magnitude!). We claim cadrille to be the first competitive multimodal CAD reconstruction approach, which can jointly handle three input modalities (point cloud, images, and text) within a unified framework.
RL for CAD and LLM
The training of modern LLMs wang2024qwen2-vl ; yang2024qwen2 is performed in several stages. The process starts with unsupervised pretraining on internet-scale data to gain broad structural knowledge of the textual domain. Then, supervised fine-tuning (SFT) is applied to train the model for specific tasks. Alignment, also known as preference optimization, helps to strengthen model’s focus on the best answers, which can be sourced from human feedback, or automatically by using elaborate RL approaches ouyang2022rlhf ; rafailov2023dpo ; shao2024deepseekmath .
Reinforcement learning has also been explored in CAD-related tasks chen2025cadcrafter ; sharma2018csgnet ; yin2025rlcad ; chao2025reinforcement . For example, RLCAD yin2025rlcad proposes an RL-based policy that interacts with a CAD engine to generate sketch-and-revolve sequences, receiving feedback based on the quality of reconstruction. Chao et. al. chao2025reinforcement applies Deep Q-Networks to recover parametric CAD models from 2D orthographic drawings in the form of loop-paths. CADCrafter chen2025cadcrafter implements a code checker to serve as an implicit reward model and fine-tunes using offline Direct Preference Optimization (DPO) rafailov2023dpo .
However, none of these works applies reinforcement learning to fine-tune large language models (LLMs) for CAD reconstruction. In this work, we are the first to explore RL fine-tuning of LLMs in this domain. We show that online RL methods like GRPO shao2024deepseekmath are particularly effective, eliminating the need for human feedback.
@@ 3 CAD Sequence Reconstruction
The task of CAD reconstruction implies recovering a CAD model given a multimodal input $q$, which can be a 3D point cloud, a set of images, or a textual description. We represent CAD models as Python scripts rukhovich2024cad-recode that, when executed, generate a parametric Boundary Representation (B-Rep) of a 3D shape. Respectively, given an input $q$, we search for a trainable policy $\pi_{\theta}$, s. t. $\pi_{\theta}(q)$ produces a token sequence $\tau$, which is essentially a text of a Python program generating a CAD model.
@@
For training a model, we derive all input modalities from ground-truth CAD models (Fig. 2). Below, we describe how each modality is constructed according to the established data generation protocols chen2025cadcrafter ; khan2024text2cad ; wu2021deepcad .
@@
Generating textual data is notably more challenging. Since our goal is accurate geometry reconstruction rather than generating a semantically relevant sample, inputs should provide detailed and comprehensive geometric information. Consequently, loose textual descriptions are generally insufficient. The necessary level of granularity is investigated in Text2CAD khan2024text2cad , where LLMs and VLMs are combined in a multi-stage sophisticated pipeline that generates textual descriptions from both the CAD sequence and rendered images.
@@ 4 Proposed Method
The architecture of our cadrille is depicted in Fig. 3. The model accepts inputs in the form of a point cloud, a set of images, or a text prompt, and outputs a Python code, that when executed, produces a CAD model. cadrille is build on top of a VLM that provides native support of text and image inputs and is already capable of generating Python code. Textual input is passed through the original embedding layer, and images are processed with an original visual encoder. The point cloud processing logic is the same as in CAD-Recode rukhovich2024cad-recode . Specifically, we use a single projection layer to embed 3D points, sample points from the surface via furthest point sampling, and do not use normals.
@@ 4.3 Limitations of SFT
Two-stage training has already been adopted in CAD-Recode rukhovich2024cad-recode , which employs supervised fine-tuning (SFT) to adapt a pretrained language model for point cloud-based CAD reconstruction. However, this strategy reveals its limitations in a cross-domain scenario: CC3D IoU is as low as 60% and the invalidity ratio (IR) is to 10%, which means that every tenth prediction fails to produce a valid output (Tab. 3, row 2). To mitigate this issue, CAD-Recode uses a test-time sampling technique. For each input query, 10 candidate Python programs are generated, and the candidate with the highest IoU is selected. After that, IoU increases to 74%, while IR drops below 0.5%. However, this improvement comes at the cost of a 10x increase in inference time. Can similar gains be achieved without sacrificing test-time efficiency?
To maintain fast and simple inference, we shift our focus to improving the training process. Training solely on procedurally generated CAD data might limit performance in real-world applications. Nevertheless, training on handcrafted models also presents challenges, e.g., prior work rukhovich2024cad-recode shows that SFT directly on the DeepCAD dataset harms performance, leading to a 10% drop in IoU.

### v3

# cadrille: Multimodal CAD Reconstruction with Reinforcement Learning

@@ Abstract
Computer-Aided Design (CAD) plays a central role in engineering and manufacturing, making it possible to create precise and editable 3D models. Using a variety of sensor or user-provided data as inputs for CAD reconstruction can democratize access to design applications. However, most existing methods focus on a single input modality: point clouds, images, or texts, which limits their generalizability and robustness, while few multimodal approaches struggle to deliver competitive quality. Leveraging advances in vision-language models (VLM), we propose cadrille, a multimodal CAD reconstruction model that takes inputs of three modalities and outputs executable Python code for CAD reconstruction. Inspired by large language model (LLM) training paradigm, we adopt a two-stage pipeline: supervised fine-tuning (SFT) on large-scale procedurally generated data, followed by reinforcement learning (RL) fine-tuning using online feedback, obtained programatically. On the DeepCAD benchmark, our SFT model outperforms existing single-modal approaches in all three input modalities simultaneously. More importantly, after RL fine-tuning, cadrille sets new state-of-the-art on as many as 10 benchmarks across three modalities and four datasets, including a real-world one. Code is avaliable at https://github.com/col14m/cadrille.
@@ 1 Introduction
Computer-Aided Design (CAD) is the core of modern engineering and manufacturing, providing the tools to create detailed and modifiable 3D models (Briere-Cote et al., 2012). Creating CAD models manually requires skills, time, and effort. To simplify this process, CAD reconstruction aims at generating CAD models directly from scanned objects, making the process faster, cheaper, and overall more accessible (Rukhovich et al., 2024).
@@
Typically, CAD models are created with a sequence of 2D sketches and 3D operations (Willis et al., 2021; Wu et al., 2021). This representation allows CAD models to be easily edited, making it prevalent in popular CAD tools like SolidWorks and AutoCAD and in CAD generation research. Most existing CAD generation methods define CAD sequences using special command tokens (Khan et al., 2024a; Wu et al., 2021). However, state-of-the-art results are obtained via mapping CAD sequences to casual Python code (Rukhovich et al., 2024). Following the same paradigm, we generate CAD models as executable Python scripts.
@@
The most well-studied input modality in CAD reconstruction is naturally a point cloud (Rukhovich et al., 2024). However, point clouds can only be obtained when a physical 3D object is available, while scanning usually requires special equipment, making the process complicated for non-experts. Images capture finer details and can be sourced using customer low-end devices (e.g., smartphone cameras), hence relaxing the hardware requirements (Chen et al., 2025; 2024). In the meantime, textual descriptions can enrich the object representation with semantic context (Khan et al., 2024b). Using various input modalities, such as multi-view images, or natural language descriptions, would make design assistance applications simple even for non-experienced users. Recent emergence of vision-language models provides a solid ground for multimodal CAD reconstruction. However, the first multimodal methods in that vein (Xu et al., 2024b; Wang et al., 2025b) are dramatically inferior to single-modal approaches, so the full potential of VLMs for CAD reconstruction is yet to be unleashed.
@@
Figure 1: Compared to state-of-the-art CAD-Recode, the only existing method that converts point clouds into Python code, cadrille has two key novelties. First, it goes beyond the standard training scheme and adapts LLM RL fine-tuning for CAD reconstruction (left). Moreover, besides point clouds only accepted by single-modal CAD-Recode, cadrille extends to images and textual descriptions, making it the first multimodal approach delivering state-of-art results (right).
@@
Existing CAD reconstruction methods face generalization issues due to how they are trained. Specifically, handcrafted CAD datasets are small and limited in diversity (Khan et al., 2024a), while models trained with procedurally generated data struggle to transfer to the real-world domain (Rukhovich et al., 2024). In the vein of the standard LLM training pipelines (Shao et al., 2024), Chen et al. (2025); Wang et al. (2025a); Guan et al. (2025) bring RL fine-tuning into the CAD reconstruction context. However, all of them perform both supervised and RL fine-tuning on the same dataset, which does not help bridging the gap between training and testing data. In contrast, we use voluminous procedurally generated data for supervised training, while valuable but scarcer handcrafted data is reserved for RL fine-tuning. This scheme eliminates the need for large-scale handcrafted data and allows the model to first generalize across the CAD domain and then specialize using preference-based objectives.
@@
Our experiments show that cadrille  outperforms existing modality-specific baselines in accuracy. Moreover, RL fine-tuning ensures validity of generated Python code, which posed a challenge for prior works. As a result, the proposed approach demonstrates impressive robustness and sets a new state-of-the-art on several CAD datasets, including a real-world CC3D (Mallis et al., 2023). Essentially, this opens up new possibilities for generalization in open-world scenarios.
@@
  • We present cadrille, an LLM-based model able to process point clouds, images, and textual inputs, and generate Python scripts for CAD reconstruction;
  • We are the first to prove that RL fine-tuning improves multimodal CAD reconstruction;
  • With a single model, we simultaneously achieve state-of-the-art results across three input modalities (point clouds, images, texts) and four datasets (DeepCAD, Fusion360, CC3D, Omni-CAD), a total of 10 benchmarks, making it the most comprehensive evaluation of CAD reconstruction methods up-to-date.
@@ 2 Related Work
Existing CAD generation methods can be classified into three categories based on CAD model representations: constructive solid geometry (CSG) (Du et al., 2018; Sharma et al., 2018; Nandi et al., 2018; Ellis et al., 2019; Tian et al., 2019; Friedrich et al., 2019; Kania et al., 2020; Ren et al., 2021; Yu et al., 2022; 2023), boundary representation (B-rep) (Wang et al., 2020; Sharma et al., 2020; Lambourne et al., 2021; Wang et al., 2022; Guo et al., 2022; Jayaraman et al., 2023; Liu et al., 2024a; b; Xu et al., 2024c; Li et al., 2025b) and CAD sequence (Wu et al., 2021; Lambourne et al., 2022; Ren et al., 2022; Xu et al., 2022; 2023; Zhang et al., 2025b; Badagabettu et al., 2024; Chen et al., 2024; Khan et al., 2024a; b; Ma et al., 2024; Mallis et al., 2024; Li et al., 2025a; Doris et al., 2025; He et al., 2025; Yuan et al., 2025; Wang et al., 2025b). In the CSG (Foley, 1996) paradigm, CAD is represented as a CSG tree constructed using boolean operations (union, subtraction, difference) of geometric primitives (e.g., cubes, cylinders, or spheres). This approach fails to express intricate shapes and is generally not well-aligned with how engineers and designers actually build CAD models. B-rep (Ansaldi et al., 1985) is a graph that describes connections between faces, edges, and vertices of a 3D model. Creating a B-Rep requires enforcing topological consistency on edges, which introduces additional complexity to the generation procedure and complicates editing of generated models.
@@
Point clouds are the most well-studied input modality in CAD reconstruction. The seminal work on point cloud-based CAD reconstruction by Wu et al. (2021) proposed encoding CAD sketch-and-extrude sequences as special tokens. Beyond that, DeepCAD, a large-scale dataset of 180k hand-crafted CAD models, was presented. Subsequent works (Dupont et al., 2024; Khan et al., 2024a; Xu et al., 2023) adopted the same CAD representation and trained on the same DeepCAD dataset. More recently, CAD-Recode (Rukhovich et al., 2024) introduced a paradigm shift by representing CAD models as Python code, providing greater expressiveness and flexibility, and released a new dataset of approx. 1 million procedurally generated CAD models.
@@
Only few recent works (Chen et al., 2024; Khan et al., 2024b; Wang et al., 2025c; Chen et al., 2025) have explored CAD reconstruction from alternative input modalities, such as single- or multi-view images and natural language descriptions. These approaches extend the DeepCAD dataset by rendering synthetic views or generating textual captions for existing CAD models. Among them, CADCrafter (Chen et al., 2025) stands out for its unified framework that handles both single- and multi-view inputs, whether rendered or real. The seminal Text2CAD (Khan et al., 2024b) uses a vision-language model (VLM) to generate detailed captions for DeepCAD shapes and then trains a model to predict CAD sequences from those textual descriptions. Its recent follow-ups (Xie & Ju, 2025; Govindarajan et al., 2025; Guan et al., 2025; Wang et al., 2025a) adapted large language models for text-based CAD reconstruction.
@@
Generally, state-of-the-art CAD reconstruction approaches are tailored to process specific input modalities with distinct architectures, while multimodal CAD reconstruction remains relatively underexplored. Recent CAD-GPT (Wang et al., 2025b) predicts a CAD model given a single image and textual description, while CAD-MLLM (Xu et al., 2024b) pioneers three-modal CAD reconstruction, yet both these methods fall behind single-modal state-of-the-art results (Rukhovich et al., 2024; Khan et al., 2024b; Chen et al., 2025) by a large margin. This makes our cadrille the first multimodal CAD reconstruction approach handling point clouds, images, and texts within a unified framework, that outperforms single-modal top-performing methods.
RL for CAD Reconstruction
Reinforcement learning is used for CAD reconstruction from images (Sharma et al., 2018; Chen et al., 2025; Zhang et al., 2025a), and from B-Rep (Yin et al., 2025). Recent LLM-based CADFusion (Wang et al., 2025a) and CAD-Coder (Guan et al., 2025) address CAD reconstruction from texts, both performing supervised and RL fine-tuning on the same DeepCAD dataset. On the contrary, we investigate RL fine-tuning for multimodal CAD reconstruction and improve the reconstruction quality by using large-scale procedurally generated data for SFT.
@@ 3 CAD Sequence Reconstruction
The task of CAD reconstruction implies recovering a CAD model given a multimodal input $q$, which can be a 3D point cloud, a set of images, or a textual description. We represent CAD models as Python scripts (Rukhovich et al., 2024) that, when executed, generate a parametric Boundary Representation (B-Rep) of a 3D shape. Respectively, given an input $q$, we search for a trainable policy $\pi_{\theta}$, s. t. $\pi_{\theta}(q)$ produces a token sequence $\tau$, which is essentially a text of a Python program generating a CAD model.
@@
For training a model, we derive all input modalities from ground-truth CAD models (Fig. 2). Below, we describe how each modality is constructed according to the established data generation protocols (Chen et al., 2025; Khan et al., 2024b; Wu et al., 2021).
@@
Generating textual data is notably more challenging. Since our goal is accurate geometry reconstruction rather than generating a semantically relevant sample, inputs should provide detailed and comprehensive geometric information. Consequently, loose textual descriptions are generally insufficient. The necessary level of granularity is investigated in Text2CAD (Khan et al., 2024b), where LLMs and VLMs are combined in a multi-stage sophisticated pipeline that generates textual descriptions from both the CAD sequence and rendered images.
@@ 4 Proposed Method
The cadrille architecture is depicted in Fig. 3. The model accepts inputs in the form of a point cloud, a set of images, or a text prompt, and outputs a Python code, that, when executed, produces a CAD model. cadrille is build on top of a VLM that natively supports text and image inputs and is already capable of generating Python code. Textual input is passed through the original embedding layer, and images are processed with an original visual encoder. The point cloud processing logic is the same as in CAD-Recode. Specifically, we use a single projection layer to embed 3D points, sample points from the surface via furthest point sampling, and do not use normals.
@@ 4.3 Limitations of SFT
Two-stage training has already been adopted in CAD-Recode, which employs supervised fine-tuning (SFT) to adapt a pretrained language model for point cloud-based CAD reconstruction. However, this strategy reveals its limitations in a cross-domain scenario: CC3D IoU is as low as 60% and the invalidity ratio (IR) almost reaches 10%, which means that every tenth prediction fails to produce a valid output (Tab. 3, row 2). To mitigate this issue, CAD-Recode uses a test-time sampling technique. For each input query, 10 candidate Python programs are generated, and the candidate with the highest IoU is selected. After that, IoU increases to 74%, while IR drops below 0.5%. However, this improvement comes at the cost of a 10x increase in inference time. Can similar gains be achieved without sacrificing test-time efficiency?
To maintain fast and simple inference, we shift our focus to improving the training process. Training solely on procedurally generated CAD data might limit performance in real-world applications. Nevertheless, training on handcrafted models also presents challenges, e.g., Rukhovich et al. (2024) shows that SFT directly on the DeepCAD dataset harms performance, leading to a 10% drop in IoU.